OpenStudy (anonymous):

Plz help.chose the point - slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of 1/2.

hartnn (hartnn):

use the equation, y-y1 = m (x-x1) where m= slope = 1/2 x1 =6 y1=-3